WI CO2 laser noise

Go down

WI CO2 laser noise

Post  Alessio on Thu Mar 12, 2009 6:48 pm

Since the WI laser is injecting noise in the dark fringe (see Virgo Logbook entry 22321), today we investigated this issue by replacing the power meter with a photodiode.
First of all, we noticed that the controller is not working properly: even if the knob is competely turned to the maximum, the laser is still working in pulsed mode instead of continuous mode as it should be. This is very clear in the first plot where the laser RIN spectrum is compared in two situations: with the controller knob to the maximum (blue curve) and with no controller at all (red curve). Anyway it is possible to see that the noisy line is still there.



We understood that this line is given by the RF driver: this driver is actually made of two separate RF drivers, this line is the beating between the two RF frequencies. For example for the laser on the North bench, this beating frequency is around 580Hz. This fact is evident by looking at figure 2, where RIN is plotted when the laser is operated with its own RF driver (red curve) or with the RF driver borrowed from the North TCS bench (blue). So changing the RF driver removes the noisy line.
The plan is to replace (next monday) the current RF driver with the spare one (now in Tor Vergata) that does not have any line in its spectrum.

avatar
Alessio

Posts : 127
Join date : 2008-11-07
Age : 46
Location : Rome

View user profile

Back to top Go down

Back to top

- Similar topics

 
Permissions in this forum:
You cannot reply to topics in this forum